OpenStudy (anonymous):

Is the equation f(x)=3x+2 a function??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes. All polynomials are functions, and the domain doesn't even have any restrictions. You can check by graphing it and using the vertical line test.

OpenStudy (anonymous):

Alright and can you tell me the inverse of that function please?!

OpenStudy (anonymous):

Four steps: 1) Let f(x) = y = 3x+2. 2) f^-1(x) is implied by x=3y+2. (Switch all the x's and y's). 3) Solve for y, in this case, y =x/3-2/3 4) y=x/3-2/3=f^-1(x).